OpenStudy (anonymous):

A box has 1 red marble, 1 blue marble, and 4 green marbles of the same shape and size. Dan draws a marble randomly from the box, replaces it, and then draws another marble randomly. What is the probability of drawing two green marbles in a row? Explain your answer.

OpenStudy (anonymous):

4/6*4/6 = 16/36 = 4/9?

OpenStudy (mathstudent55):

The probability of an event happening is: \(p(event) = \dfrac{number ~of~desired~outcomes}{total~ number ~of~ outcomes}\) If two events are independent of each other, the probability of both events happening is the product of their individual probabilities.

OpenStudy (mathstudent55):

You are correct. \(p(green~ marble) = \dfrac{4}{6} \) \(p(green~marble ~followed~by ~green~marble) = \dfrac{4}{6} \times \dfrac{4}{6} = \dfrac{4}{9} \)

OpenStudy (anonymous):

I am trying to put it in worded form because it value 10pts